OpenStudy (anonymous):

Calculate the 100th partial sum of the sequence. 1, 3, 7, 10, ... Do i just add all one hundred of these numbers!? 1+3+7+10+13 etc!?!?

OpenStudy (welshfella):

that would be pretty tedious!

OpenStudy (welshfella):

though I cant see a pattern in this series you sure you have the question right?

OpenStudy (anonymous):

yea lol im wondering that too

OpenStudy (ahsome):

What is the pattern of the numbes?

OpenStudy (jdoe0001):

can you post a quick screenshot of the material?

OpenStudy (bibby):

I think he just made it up for an example

OpenStudy (welshfella):

are you sure the second term is not 4?

OpenStudy (ahsome):

I see. Yes, you would have to manually solve it, or display it in summation

OpenStudy (welshfella):

that would make it an arithmetic sequence

OpenStudy (anonymous):

NO WAIT CRAM CRABBIT DAG GUMMIT I MADE A MISTAKE

OpenStudy (anonymous):

The Sequence is 1, 4, 7, 10, i'm so sorry

OpenStudy (welshfella):

???? lol

OpenStudy (anonymous):

I'm gonna go nuts i need to CEASE with these flipping sequences

OpenStudy (jdoe0001):

hmm can you post a quiick screenshot of the material? thus we could see what might be

OpenStudy (welshfella):

thats an arithmetic sequence with commin difference of 3

OpenStudy (anonymous):

yeap x-x

OpenStudy (jdoe0001):

well 1 + 3 = 4 though 2nd term is 3 there

OpenStudy (welshfella):

do you know the formula for the sum of n terms?

OpenStudy (jdoe0001):

ohhh.. hmmm

OpenStudy (welshfella):

@jdoe0001 - he later corrected it to 1,4,7,10

OpenStudy (anonymous):

I'm trying to crack at it, but I can't think of a single formula. I could pop those numbers into a calculator and get an f(x) formula easy, but not for a blippin recursive sequence, god i hate these things

OpenStudy (welshfella):

Sn = (n/2) [2a1 + (n - 1)d] where a1 = first term and d = common difference

OpenStudy (welshfella):

here a1 = 1, n = 100 and d = 3

OpenStudy (jdoe0001):

\(\large { \textit{what's }a_{\color{brown}{ 100}}?\qquad well \\ \quad \\ a_{\color{brown}{ n}}=a_1+({\color{brown}{ n}}-1)d\qquad d\to \textit{common difference}\to 3 \\ \quad \\ a_{\color{brown}{ 100}}=a_1+({\color{brown}{ 100}}-1)\cdot 3 \\ \quad \\ \\ \quad \\ S_{\color{brown}{ n}}=\cfrac{{\color{brown}{ n}}}{2}(a_1+a_{\color{brown}{ n}})\impliedby \textit{sum of an arithmetic sequence} \\ \quad \\ {\color{brown}{ n}}=100\qquad S_{\color{brown}{ 100}}=\cfrac{{\color{brown}{ 100}}}{2}(a_1+a_{\color{brown}{ 100}}) }\)

OpenStudy (anonymous):

So 397!?

OpenStudy (anonymous):

wait what has jdoe got here

OpenStudy (welshfella):

nop try again

OpenStudy (jdoe0001):

eheheh

OpenStudy (anonymous):

oh cram it i multiplied something instead of divided gAK

OpenStudy (ahsome):

You could also go solve the summation for it\[\Huge{\sum\limits_{i=0}^{99} (3i+1)} \]

OpenStudy (welshfella):

S100 = 50[2 + (100-1)3]

OpenStudy (anonymous):

oh how i prefer sigma equations....

OpenStudy (anonymous):

Alright i'm clearly getting something wrong because i'm still getting 397

OpenStudy (anonymous):

oh WAIT

OpenStudy (anonymous):

wait no yeah i'm going something wrong

OpenStudy (welshfella):

work out inner bracket first 99*3 then add 2 then multiply by 50

OpenStudy (anonymous):

it was the parentheses, dag gummit, so do you have 14950?

OpenStudy (welshfella):

yea

OpenStudy (anonymous):

fantastic

OpenStudy (welshfella):

remember PEDMAS